OpenStudy (anonymous):

How do I divide x^8+x^4/x? I keep getting x^8+x^3 and I do not think that is right...is that right? ugh thanks in advance

OpenStudy (jhannybean):

look back at the previous question you posted,i solved it there.

OpenStudy (jhannybean):

:\

OpenStudy (johnweldon1993):

the +x^3 part is right....but what do you multiply 'x' by to get x^8?

OpenStudy (anonymous):

ugh, I have this all over the place..I have no idea.

OpenStudy (jhannybean):

\[\large \frac{x^8+x^4}{x}\]\[\large \frac{x^8}{x} + \frac{x^4}{x}\] just solve each one separately

OpenStudy (jhannybean):

so for the first part you'd have \[\large x^{(8-1)} + x^{(4-1)}\]

OpenStudy (jhannybean):

i mean first and second* and @johnweldon1993 can continue....sorry john :|

OpenStudy (anonymous):

10x?

OpenStudy (johnweldon1993):

no, you already did it before....go ahead and finish it @Jhannybean :)

OpenStudy (anonymous):

I'm lost

OpenStudy (jhannybean):

I don't think you can add exponents when you're adding variables... :;\ only when multiplying.

OpenStudy (anonymous):

oh damn

OpenStudy (jhannybean):

what would you get when you simplified \[\large x^{(8-1)} + x^{(4-1)}\] though?

OpenStudy (anonymous):

x^7 + x^3

OpenStudy (anonymous):

that is not what I get but, I think that;s it

OpenStudy (jhannybean):

okay, it can go a little further

OpenStudy (jhannybean):

See how you have a common base variable "x" ?

OpenStudy (jhannybean):

even though the two adding have DIFFERENT powers,we can FACTOR out a GFC, which in this case is x^3 :)

OpenStudy (anonymous):

you guys rock! thanks

OpenStudy (jhannybean):

so you would have \[\large x^3(x^4+1)\]

OpenStudy (anonymous):

I have not had math like this in like 20 years - still a work in prgress

OpenStudy (jhannybean):

Ohh haha yeah it happens. always have to keep practicing!

OpenStudy (anonymous):

apparently...thanks again :)

OpenStudy (jhannybean):

noo problem!